With the problem: 4x/5+3/2=2x the answer is x=5/4
How would the equation be to check my answer?
What is the step by step process to solve 4(5/4)/5+3/2=2(5/4)?

Answer by rfer(16322)

Your setup is great, just finish it.
(20/4)/5+3/2=10/4
5/5+3/2=10/4
1+3/2=10/4
2/2+3/2=10/4
5/2=10/4
10/4=10/4